The Dawson Company has just completed a number of budgets for the coming year. The cost of goods manufactured schedule, the pro forma income statement, and the balance sheet still have to be completed.

The following information is available as 12/31/X8.

Prior Year Balance Sheet

Assets
Cash $45,000
Accounts Receivable $55,000
Materials Inventory $40,000
Work-in-Process Inventory $30,000
Finished Goods Inventory $36,000
Prepaid Expenses $20,000
Plant and Equipment $500,000
Accumulated Depreciation ($140,000)
Other Assets $22,000
Total Assets $608,000

Liabilities and Equity

Accounts Payable $103,000
Other Current Liabilities $42,000
Income Taxes Payable $25,000
Long-Term Debt $300,000
Total Liabilities $470,000
Common Stock $100,000
Retained Earnings $38,000
Total Equity $138,000
Total Liabilities and Equity $608,000

Information From Recent Budgets for the Coming Year

  1. Projected sales are $2,080,000 (13,000 units).
  2. Projected direct materials purchases are $525,000.
  3. Projected direct materials usage is $510,000.
  4. Projected direct labor expense is $420,000.
  5. Projected overhead is $390,000.
  6. Projected selling expenses are $130,000.
  7. Projected administrative expenses are $310,000.
  8. Projected cash collections are $1,805,000.
  9. Projected payments for materials (accounts payable) are $550,000.
  10. Projected payments for other operating expenses (other current liabilities) are $1,155,000.
  11. Projected depreciation expense is $60,000 and is already included in manufacturing overhead.

Additional Information That Is Available

  1. The expected tax rate is 20%.
  2. The company is planning a stock issue of $50,000.
  3. Income taxes are paid 3 months after year-end.
  4. The company anticipates purchasing a new patent for $20,000 during the year.
  5. Work-in-Process Inventory is expected to decrease by $2,500.
  6. Finished Goods Inventory is expected to increase by $9,000.
  7. Due to insurance rate increases, it is expected that prepaid expenses will increase by $5,000.

Investment Information

  1. A purchase of additional equipment for $80,000 is expected on January 2, 20X9.
  2. The purchase will be made using $50,000 cash and long-term debt will be increased by $30,000.

Long-Term Debt Information

  1. All long-term debt will have a 9% annual rate.
  2. A payment of $50,000 including both principal and interest will be made on December 31, 20X9.
Answer #1

Cost of Goods Manufactured,

Add Particuars Direct Material Used Direct Labour Used Manufacturing OH Depreciation (factory Plant) Add Add Amount $ 5,10,00

Proforma Income Statement,

Proforma Income statement for the period 01/01/20x9 to 31/12/20X9 Particulars Amount $ Revenue 20,80,000 Cost of goods sold 5

Balance sheet Adjustment 20X8 20x9 18,55,000 20,80,000 5,25,000 18,55,000 18,05,000 5,10,000 2,500 Assets Cash Accounts Recei

Workings,

Closing Stock of Inventory Particular Material inventory Add: Purchase Less: Usage Closing Stock of Inventory Amount % 40,000

Interest on Bond issue - Op $ 300000 Plus issued in 2nd Jan,20X8 $ 30000 = 330000 * 9% = $ 29700 interest

Payment of Bond = 50000-29700 = 20300 Principal component

Cash Opening Balance Sales Realisation Share issued 45,000 18,05,000 50,000 19,00,000 Less: Accounts Payable Other Current Li

Paticulars Manpower Expenses Other Overhead Selling Expenses Administrative expenses Total other Current Liability Payable Am